If you invest $10,000 today and it grows at annual rate of 15% (compounded monthly), how many months it will take to grow to $20,000?
Answer #1

use financial calculator as

FV=20000

PV=-10000

I/Y=15/12=1.25

PMT=0

Click CPT

Click N=55.80 months

the above is answer..